OpenStudy (anonymous):

Find the rate of change between the two points: (6, 8) and (5, 6) where x is in days and y is in pounds.

OpenStudy (anonymous):

difference of y / difference x x: 5-6=-1 y 6-8 =-2 -1 days / -2 pounds = 1/2 days per pound